14. Write an equation of the line with slope -2 and passes through the point (5.-7).

can u solve it asap please with explain
Answer #1

Equation of line point (5,-7) and (1,, 41) passing through slope -2 is (m) y-y = m(2-x) Y-(-) = -2(2-5) Y+7= -2x +10 y = -2x+